# Solving the Traveling Salesman Problem via Different Quantum Computing Architectures

### v1

@@ Abstract
QAOA and QPE were tested on the IBM Quantum platform. The QUBO-Ising method was explored using the D-Wave quantum annealer, which operates on superconducting Josephson junctions, and the QCI Dirac machine, a nonlinear optoelectronic Ising machine. Gate-based quantum computers demonstrated accurate results for small TSP instances in simulation. However, real quantum devices are hindered by noise and limited scalability. Circuit complexity grows with problem size, restricting performance to TSP instances with a maximum of 6 nodes.
In contrast, Ising-based architectures show improved scalability for larger problem sizes. SQUID-based Ising machines can handle TSP instances with up to 12 nodes, while nonlinear optoelectronic Ising machines extend this capability to 18 nodes. Nevertheless, the solutions tend to be suboptimal due to hardware limitations and challenges in achieving ground state convergence as the problem size increases. Despite these limitations, Ising machines demonstrate significant time advantages over classical methods, making them a promising candidate for solving larger-scale TSPs efficiently.
@@ III-D1 Hamiltonian formulation on the QCI Dirac platforms
The QCI dirac machines are photonic quantum computers that can solve BQM, QCBO and QUBO problems in a similar way that the D-wave quantum annealers can solve the problems. The have been shown to additionaly solve integer linear programming models and mixed integer linear programming models with a solution space spanning a range of integers without explicitly converting such a problem to binary. This allows us to use ILP formulations directly without any conversions to define the TSP to be solved on the QCI platforms. Two most popular ILP formulations that are being tested here are the Miller–Tucker–Zemlin [26] formulation and the Dantzig–Fulkerson–Johnson formulation[27].
@@ III-D4 Hardware architecture
The Dirac1 machine is an FPGA based Optical Parametric Oscillator(OPO), that uses a non-linear crystal such as Periodically Poled Lithium Niobate (PPLN). Optical Parametric Oscillators (OPOs) are devices that generate new frequencies by mixing two light waves in a nonlinear crystal. The PPLN is a highly efficient medium for these nonlinear wavelength conversion processes. It is used for frequency doubling, difference frequency generation, sum frequency generation, optical parametric oscillation, and other nonlinear processes [29].
@@
Since no explicit embedding step is required as opposed to the D-wave platform [22], the QCI Dirac machines are able to solve TSP instances of sizes beyond 12 nodes.
@@ III-D5 Solution for an 18 node TSP
The maximum size of the TSP that can be solved reliably without violating the constraints is currently 18 nodes. Figures 20 is a demonstration of the QCI dirac machine solving a TSP instance of 18 nodes.

### v2

@@ Abstract
QAOA and QPE were tested on the IBM Quantum platform. The QUBO-Ising method was explored using the D-Wave quantum annealer, which operates on superconducting Josephson junctions, and the Quantum Computing Inc (QCi) Dirac-1 entropy quantum optimization machine. Gate-based quantum computers demonstrated accurate results for small TSP instances in simulation. However, real quantum devices are hindered by noise and limited scalability. Circuit complexity grows with problem size, restricting performance to TSP instances with a maximum of 6 nodes.
In contrast, Ising-based architectures show improved scalability for larger problem sizes. SQUID-based Ising machines can handle TSP instances with up to 12 nodes, while entropy computing implemented in hybrid optoelectronic components extend this capability to 18 nodes. Nevertheless, the solutions tend to be suboptimal due to hardware limitations and challenges in achieving ground state convergence as the problem size increases. Despite these limitations, Ising machines demonstrate significant time advantages over classical methods, making them a promising candidate for solving larger-scale TSPs efficiently.
@@ III-D1 Hamiltonian formulation on the QCi Dirac platforms
QCi is currently offering hybrid quantum optimization machines that can solve discrete and quasi-continuous variables (Dirac-3) or QUBO/Ising problems (Dirac-1) like D-wave quantum annealers. Dirac-3 has been shown to additionaly solve integer linear programming models and mixed integer linear programming models with a solution space spanning a range of integers without explicitly converting such a problem to binary. This allows us to use ILP formulations directly without any conversions to define the TSP to be solved on the QCI platforms. Two most popular ILP formulations that are being tested here are the Miller–Tucker–Zemlin [26] formulation and the Dantzig–Fulkerson–Johnson formulation [27].
@@ III-D4 Hardware architecture
Entropy quantum computing was recently introduced as an efficient optimization technique that leverages the principles of entropy minimization to guide quantum state evolution toward optimal solutions of a Hamiltonian [33]. In a photonic realization, qudits are encoded as a superposition of photon numbers in time bins that evolve in an optical fiber loop which embeds a desired Hamiltonian as a dissipative operator. This effectively emulates imaginary time evolution, in which the propagation of higher-energy eigenstates is subject to dissipation and decoherence while the lower-energy eigenstates are promoted in this evolution.
The current generation of the Entropy Quantum Computing (EQC) hardware at QCi, named Dirac-3, is a hybrid system that harnesses the high-speed parallel processing capabilities of photonics and the quantum nature of light with the precise control and programmability of electronic circuits. By leveraging photonic systems for entropy-driven state evolution—enabled by their natural ability to manipulate complex optical fields—and utilizing electronics for state initialization, feedback, and fine-tuning, hybrid entropy computing achieves efficient optimization. Such systems can explore large solution spaces rapidly through photonic modes propagation while employing electronic components to refine solutions iteratively. This synergy offers a pathway for solving challenging optimization problems in areas such as machine learning, signal processing, and network design, with the potential for high speed, low power consumption, and scalability.
The EQC hardware Dirac-1 targets optimization of a polynomial objective function $E$ in the following form:
$E=\sum_{i}C_{i}x_{i}+\sum_{i,j}J_{ij}x_{i}x_{j}.$ (51)
where, $x_{i}\in\{0,1\}$ are optimization variables, $C_{i}$ are real-valued coefficients of linear terms, and $J_{ij}$ represent two-body interaction coefficients that are real numbers subject to the tensor $J$ being symmetric under all permutations of the indices. Governed by the physical mechanism of operation of the EQC hardware, $x_{i}$ represents photon numbers in the $i$’th time bin degrees of freedom.
A major advantage of EQC compared to other quantum annealers is its flexible variable encoding capability [33]. While, in the form discussed above, EQC allows for solving continuous-variable optimization problems, by allocating multiple photon time bins to one variable, one can effectively encode binary or integer variables for solving combinatorial optimization problems in the forms of binary, integer or even mixed-integer problems.
@@
With the advantage of all-to-all connectivity no explicit embedding step is required as opposed to the D-wave platform [22], the QCI Dirac machines are able to solve TSP instances of sizes beyond 12 nodes.
@@ III-D5 Solution for an 18 node TSP
The maximum size of the TSP that can be solved reliably without violating the constraints is currently 18 nodes. Figures 20 is a demonstration of the Dirac-1 machine solving a TSP instance of 18 nodes.